OpenStudy (anonymous):

Solve each system by elimination. 28-4(y+1)=3(x-y)+4 -5(x+4)-y+3=28-5(2x+5)

OpenStudy (anonymous):

@ashrafh

OpenStudy (anonymous):

z= 12+28 = 40

OpenStudy (anonymous):

replace z with x

OpenStudy (anonymous):

\[28-4y-4=3x-3y+4\] \[20-y=3x\] \[20-3x=y\] put this into the second equation

OpenStudy (anonymous):

it will give yoeverything in terms of x